The weight of water provided the force on the wheel to turn the axle. If the axle provides 3750N of force, and the effective weight of the water can provide a tangent force on the wheel of 325N, what is the minimum radius of the wheel if the axle radius is 55 cm.
1
Expert's answer
2011-02-03T08:09:35-0500
Let's write the condition of equilibruim of the torques:
F1d1 = F2d2
3750*0.55 = 325*d2
d2 = 3750*0.55/325 ≈ 6.35 m.
Answer: the minimum radius of the wheel is equal approximately to 6.35 m.